OpenStudy (anonymous):

what is the equationof the line that passes though (4,-1) and (-2,3)

OpenStudy (cwrw238):

first find the slope of the line using the formula y2 - y1 ------ x2 - x1 where the 2 points are (x1,y1) and (x2, y2)

OpenStudy (cwrw238):

in you case (x2,y2) can be (-2,3) in other words x2 = -2 and y2 = 3

OpenStudy (anonymous):

you can assign point P1 (4, -1) and point P2 (-2, 3)... as @cwrw238 said, find the slope first... then you can complete the line equation using the Point-Slope form... \[(y - y1) = m (x - x1) \]

OpenStudy (cwrw238):

m is the slope and you are given x1 and y1

OpenStudy (anonymous):

x+2=-2/3(y-4)